OpenStudy (anonymous):

4sin 60degrees -3tan 135degrees. Find the exact value

OpenStudy (anonymous):

\[3+2 \sqrt{3} \]

OpenStudy (anonymous):

Well I know the answer, but I don't know how to get the answer.

OpenStudy (anonymous):

\[\text{Sin}[60{}^{\circ}]=\frac{\sqrt{3}}{2},\text{Tan}[135{}^{\circ}]=-1,4\frac{\sqrt{3}}{2}-3(-1)=3+2 \sqrt{3} \]

OpenStudy (anonymous):

Thank you so much!